# Removing constraints of 4D-STEM with a framework for event-driven acquisition and processing

### v1

@@ Introduction
In terms of read-out mechanism, most HPAD like e.g. those based on the Medipix3 chip [4, 38] are still very similar to the CCD and CMOS detectors that have been used routinely in TEM. This read-out mechanism is typically called frame-based, where a value for every pixel is recorded after a chosen exposure period. The maximum frame rate of current generation frame-based HPAD is in the order of 100 kHz [11], making them still at least an order of magnitude slower than what is frequently used for single-pixel detectors. When no fast electrostatic beam blanker is available (synchronized with the scanning), this minimum frame time puts a lower limit on the electron dose for a scan with a workable fixed current and number of scan positions.
@@
The cryo-EM and microED community have also realized workflows exploiting the sparsity of electron detection, for example using a Timepix3-based DED [51, 52] or using the electron event representation (EER) as an in-line compression method [53, 54, 55]. In the case of the EER, the readout mechanism of the detector is not directly event-driven like in the case of Timepix3 and no TOA is available. Instead, event extraction from raw frames can be performed on a hardware level before crossing the bandwidth-limited communication channel to the user’s PC. Localization of the electron impact with sub-pixel precision can be performed from the charge distribution if the frame is sufficiently sparse and coincidence losses can be avoided. EER is implemented on for example the Thermo Fisher Scientific Falcon 4i DED and the Direct Electron Apollo DED [56].
@@
An important consideration for event-driven operation is the difference between events per second and electrons per second, because of the activation of a cluster of neighboring pixels. For a frame-based detector, the cluster effect does not increase the data rate, and only affects the modulation transfer function (MTF). For an event-driven DED, the cluster effect increases the data rate and, as a result, it depends also on the chip’s discriminator threshold. However, when TOA information is available, this can be corrected via a so-called declustering process, as has been demonstrated for sub-pixel precision imaging [51, 59]. For high-resolution 4D-STEM imaging, it was reported not to have a very significant effect on the reconstructed image [49]. It was also reported that this de-clustering can be performed in real-time [60], which could provide a data rate reduction to the further processing pipeline. Histograms of the cluster size distribution for the experimental datasets used in this work are provided in the supplementary information. For the theoretical assessments made in the rest of this work, an average cluster size of 3 will be assumed, which is a typical value for a 200 kV electron at a suitable threshold and sensor layer thickness [49, 59].
@@
Noteworthily, the event-driven setup has mostly been demonstrated for high-resolution STEM imaging. To demonstrate that these alternative scaling laws can be beneficial in a broad range of applications, figure 3 shows an example of a nano-beam 4D-STEM dataset of a commercial ZIF-8 metal-organic framework using 30 scans of 2048$\times$2048 positions using 5 $\mu$s dwell time, covering a field-of-view of approximately 10 $\mu$m. Thanks to the event representation, the total data size is approximately 34 GB, while an uncompressed 8 bit frame format would require approximately 33 TB. Further experimental details are provided in the supplementary information. Figure 3 demonstrates how, thanks to the event mode of the detector, the resolution obtained in real and reciprocal space simultaneously is much closer to being directly limited by physics rather than practical constraints, while maintaining high speed and, as a result, low electron dose. Since the data size only scales with the number of counts, it does not depend on the fractionation of a measurement into multiple scans. For a frame-based detector, the splitting of a single scan with dwell time $\tau$ into $n$ faster scans using $\frac{\tau}{n}$, which in the first place is bound to the lower limit on the frame time, comes with an increase of the data size of a factor $n$. The ability to split a measurement into multiple high-speed measurements brings insight into the dynamics of the sample, that would otherwise be hard to achieve, at no cost. Furthermore, it has been shown that for certain samples, a reduction of the dose rate, which can be achieved through such a fractionation into multiple scans, can significantly reduce the induced electron beam damage [66, 67]. Additionally, the fractionation of a scan into multiple fast scans also allows to correct sample drift [68].
Figure 2: a) Data size of a 4D-STEM dataset as a function of number of probe positions, for different acquisition parameters, in the event-driven and frame-based scenarios. b) Data size ratio of a 64 bit event representation and an 8 bit frame representation for a detector of 512$\times$512 pixels, as a function of beam current and illumination time. c) Number of operations required to calculate a virtual detector count and CoM value for a single probe position, in the event-driven and frame-based scenarios. For all, an average cluster size of 3 is assumed. It must be noted that not all scenarios considered here can actually be realized with current generation HPAD, such as a beam current of 50 pA with event-driven detectors or the fast frame rates with frame-based detectors.
@@
This need for live processing of large datasets generated by 4D-STEM has created a trend toward the incorporation of high-performance computing (HPC) infrastructure into EM workflows [78, 79, 80, 12, 81, 82]. Although in certain cases the need for such infrastructure might be inevitable, this work shows that, in many cases, efficient representation of electron data enables in-line processing using standard microscope control hardware—such as a desktop computer—with sufficient performance. HPC workflows do not only come with large costs and limited accessibility, but they potentially also have higher latency than local computing infrastructure. Such high latency is an important limitation for HPC when aiming for truly live processing, for example in using reconstructions from data streams in a search mode when finding a sample or optimizing the optics of the microscope. This is an important consideration as, for example, the benefits provided by dose-efficient 4D-STEM methods are strongly mitigated when the user is still limited to less appropriate methods like ADF imaging when setting up for acquisition, and is forced to choose between either damaging the sample anyway or working blindly. These considerations regarding the implications of HPC for 4D-STEM data processing, like cost, barrier to entry, accessibility and latency, are important to consider now, as they determine the future of these 4D-STEM methods and their integration in the wider field. Aiming for accessible, fast and efficient workflows increases the chances of 4D-STEM methods to establish themselves as routine and widespread tools for the investigation of e.g. life science specimens or other beam-sensitive materials.
@@
In the discussion above, the distinction between frame-based and event-driven detectors is made on the level of the chip application-specific integrated circuit (ASIC), like in the case of the Medipix3 and Timepix3 chips. However, the read-out mechanism of a DED does not necessarily have to be considered fixed to that of the ASIC, as it is possible to transform the initial data stream between different representations [53, 54, 83, 12]. This transformation could happen on a low hardware level, before the data stream reaches the user’s processing computer, for example by performing run-length encoding using a field-programmable gate array (FPGA). Many of the properties of the DED will remain determined by the configuration of the ASIC, like the maximum beam current, the maximum speed, and the information content, e.g. the presence of TOA information. Consequently, this will anyhow determine the range of applications allowed by the DED, such as e.g. time-resolved measurements [46, 47, 43, 45, 44, 48]. Still, the initial configuration of the DED is not capital for the arguments described above, as these simply consider sparse representations of electron data rather than the more classical dense representation.
@@
Currently, the event-driven DED for which acquisition control is already implemented include the Amsterdam Scientific Instruments CheeTah T3 and a custom prototype built from a commercial Advacam AdvaPIX T3. The frame-based Quantum Detectors MerlinEM [91, 92] is also supported, which offers acquisition in a 1 bit mode, having interests for straightforward sparsification [40]. Regarding scan control, the framework currently supports the use of the Quantum Detectors scan engine, which provides possibilities for various triggering and synchronization functionalities including e.g. line triggering, pixel triggering, master clock generation and pattern scanning. These functionalities are aligned with the processing pipeline to enable seamless synchronization in many configurations. For the integration of the microscope column’s hardware and functionalities, there currently is support for Thermo Fisher Scientific instruments employing the temscript Python wrapper [93]. Integration of JEOL systems is planned for future work. A graphical user interface (GUI) is available that provides centralized and synchronized control over all hardware components. As such, live feedback over experimental output is accessible. Thanks to this centralized control over all components, advanced measurements schemes are available for routine use, including automated 4D-STEM tomography, alternative scanning 4D-STEM and live ptychography.
@@
The live image reconstruction pipeline is implemented in the C++ programming language. The rather high performance requirement for the processing of the raw data streams provided by the detector, encompassing the decoding of raw events to usable electron coordinates, warrants the need for an efficient implementation in a high-performance language. In the present case, it is fulfilled by the C++ language. Another framework for the processing of event-driven STEM experiments, dedicated towards EELS, was developed by Auad et al. [50] and is based on the Rust programming language.
@@
In the case of the event-driven kernel-additive approach, the computational complexity of the reconstruction follows $\text{O}(N\cdot R^{2})$, with $N$ the number of events and $R^{2}$ the total amount of pixels in the kernel. As such, the number of floating point operations per second (FLOPS) that has to be made to live process a complete event-driven data stream, in the GPRI framework, can be directly estimated from the beam current and kernel size. A single-chip Timepix3 detector may provide a maximum of about 80 million events per second, which corresponds to a beam current of a few pA depending on the cluster size. A realistic size for a GPRI kernel is 64 $\times$ 64 pixels in the reconstruction window, and a single-precision floating-point representation is sufficient. In this scenario of a single precision 64 $\times$ 64 kernel, and assuming a beam current of 5 pA with an average cluster size of 3, the number of additions to be performed per second is just under 400 GFLOP. Going further, a more detailed evaluation of the required FLOPS, as a function of beam current or frame rate and kernel size, is presented in figure 5. Standard graphical processing units (GPU) available in desktop computers, like the Nvidia RTX series, have a theoretical peak performance in the tens to hundreds of TFLOPS regime. While the FLOPS performance achieved in practice is quite significantly less than this theoretical value, it then remains clear that the computational requirements are within an achievable order of magnitude. More generally, this means that, in the GPRI framework, there is no requirement for advanced hardware, that would go beyond standard consumer hardware, to meet the computational requirements of live analytical ptychography reconstructions from DED data streams. This nevertheless still requires a performant implementation of the processing pipeline, that can sufficiently leverage the parallel nature of a GPU.
Figure 5: Assessment of the computational requirements of a kernel-additive method like GPRI as a function of kernel size and data rate. In a, the event hit rate and associated beam current (with cluster size 3) is used, showing the maximum hit rate of current generation event-driven DED. In b, the frame rate and associated frame time is used, situating current generation DED [9, 91, 11]. To enable the comparison in b, a detector of 128 $\times$ 128 pixels is assumed, which is not the actual total number of pixels for the detectors marked with a star. The required FLOPS range from the order of GFLOPS to 100 TFLOPS, which is within the theoretical peak performance of a single current-generation consumer GPU.
@@
In this subsection, a demonstration is made of the live processing performance using a consumer desktop computer consisting of a 16 core AMD Ryzen threadripper pro 5995wx CPU and an Nvidia RTX4090 GPU. While this is a rather powerful setup, desktop-wise, it of course remains extremely modest in comparison to typical HPC infrastructure. For repeatability, the evaluations are based on raw recorded datasets stored on an NVMe solid-state disk. However, the results are also representative for raw data streams coming from a socket connected to the detector. For GPRI, GPU acceleration is achieved using a combination of Pytorch’s LibTorch API [98] and custom CUDA kernels [99]. Before making an assessment of the processing speeds of the ptychographic reconstructions using the GPRI framework, an evaluation is made of the performance of the underlying evenTem pipeline for parsing the raw packets to $\left(r_{x},r_{y},k_{x},k_{y}\right)$ coordinates. To this end, a virtual ADF reconstruction is made using a single CPU thread for reading data and a single CPU thread for parsing the events. The riCoM method is also used for comparison, as it has already been shown to provide live dose-efficient reconstructions, based directly on the event stream.
@@ Discussion
  • When the time resolution of the detector is in the nanosecond or better regime, like in the case of Timepix3 and Timepix4, the detector does not impose the practical limit on the scan speed. Instead, this limit is imposed by response time of the scan system [49, 100].
@@
Overall, this means that Timepix4-based detectors hold promises to become an ideal all-around solution for STEM, thus reinforcing the need to drive the integration of event-driven operation into electron microscopy workflows. The first experimental setups integrating the Timepix4 chip have already been realized [59], although the maximum data rate performance has not yet been reported.
From the assessments presented in this work, it is clear that the event format still provides numerous benefits over a dense format in many of the scenarios enabled by Timepix4. Furthermore, to keep up with its maximum data rate, the current processing pipeline will have to allow further parallelization. In the current workflow, a single CPU thread is used for event parsing. If events can be parsed independently of each other, an expansion to multiple CPU threads or a GPU should be straightforward. As discussed above, when synchronization is ensured by the use of a master clock, events can indeed be treated independently.

### v2

@@ Introduction
In terms of read-out mechanism, most HPAD like e.g. those based on the Medipix3 chip [4, 38] are still very similar to the CCD and CMOS detectors that have been used routinely in TEM. This read-out mechanism is typically called frame-based, where a value for every pixel is recorded after a chosen exposure period. The maximum frame rate of current generation frame-based HPAD is in the order of 100 kHz [11], making them still at least an order of magnitude slower than what is frequently used for single-pixel detectors like in ADF-STEM. When no fast electrostatic beam blanker is available (synchronized with the scanning), this minimum frame time puts a lower limit on the electron dose for a scan with a workable fixed current and number of scan positions.
@@
The cryo-EM and microED community have also realized workflows exploiting the sparsity of electron detection, for example using a Timepix3-based DED [51, 52] or using the electron event representation (EER) as an in-line compression method [53, 54, 55]. In the case of the EER, the readout mechanism of the detector is not directly event-driven as in the case of Timepix3 and no TOA is available. Instead, event extraction from raw frames can be performed on a hardware level before crossing the bandwidth-limited communication channel to the user’s PC. Localization of the electron impact with sub-pixel precision can be performed from the charge distribution if the frame is sufficiently sparse and coincidence losses can be avoided. EER is implemented on for example the Thermo Fisher Scientific Falcon 4i DED and the Direct Electron Apollo DED [56].
@@
An important consideration for event-driven operation is the difference between events per second and electrons per second, because of the activation of a cluster of neighboring pixels. For a frame-based detector, the cluster effect does not increase the data rate, and only affects the modulation transfer function (MTF). For an event-driven DED, the cluster effect increases the data rate and, as a result, it depends also on the chip’s discriminator threshold. However, when TOA information is available, this can be corrected via a so-called declustering process, as has been demonstrated for sub-pixel precision imaging [51, 59]. For high-resolution 4D-STEM imaging, it was reported not to have a very significant effect on the reconstructed image [49]. It was also reported that this de-clustering can be performed in real-time [60], which could provide a data rate reduction to the further processing pipeline. Histograms of the cluster size distribution for the experimental datasets used in this work are provided in the supplementary information. For the theoretical assessments made in the rest of this work, an average cluster size of 3 pixels will be assumed, which is a typical value for a 200 kV electron at a suitable threshold and sensor layer thickness [49, 59].
@@
It is noteworthy that the event-driven setup has mostly been demonstrated for high-resolution STEM imaging. To demonstrate that these alternative scaling laws can be beneficial in a broad range of applications, figure 3 shows an example of a nano-beam 4D-STEM dataset of a commercial ZIF-8 metal-organic framework using 30 scans of 2048$\times$2048 positions using 5 $\mu$s dwell time, covering a field-of-view of approximately 10 $\mu$m. Thanks to the event representation, the total data size is approximately 34 GB, while an uncompressed 8 bit frame format would require approximately 33 TB. Further experimental details are provided in the supplementary information. Figure 3 demonstrates how, thanks to the event mode of the detector, the resolution obtained in real and reciprocal space simultaneously is much closer to being directly limited by physics rather than practical constraints, while maintaining high speed and, as a result, low electron dose. Since the data size only scales with the number of counts, it does not depend on the fractionation of a measurement into multiple scans. For a frame-based detector, the splitting of a single scan with dwell time $\tau$ into $n$ faster scans using $\frac{\tau}{n}$, which in the first place is bound to the lower limit on the frame time, comes with an increase of the data size of a factor $n$. The ability to split a single scan into multiple high-speed scans brings insight into the dynamics of the sample, that would otherwise be hard to achieve, at no cost. Furthermore, it has been shown that for certain samples, a reduction of the dose rate, which can be achieved through such a fractionation into multiple scans, can significantly reduce the induced electron beam damage [66, 67]. Additionally, the fractionation of a scan into multiple fast scans also allows to correct sample drift [68].
Figure 2: a) Data size of a 4D-STEM dataset as a function of number of probe positions, for different acquisition parameters, in the event-driven and frame-based scenarios. b) Data size ratio of a 64 bit event representation and an 8 bit frame representation for a detector of 512$\times$512 pixels, as a function of beam current and illumination time. c) Number of operations required to calculate a virtual detector count and CoM value for a single probe position, in the event-driven and frame-based scenarios. For all, an average cluster size of 3 pixels is assumed. It must be noted that not all scenarios considered here can actually be realized with current generation HPAD, such as a beam current of 50 pA with event-driven detectors or the fast frame rates with frame-based detectors.
@@
This need for live processing of large datasets generated by 4D-STEM has created a trend toward the incorporation of high-performance computing (HPC) infrastructure into EM workflows [78, 79, 80, 12, 81, 82]. Although in certain cases the need for such infrastructure might be inevitable, this work shows that, in many cases, efficient representation of electron data enables in-line processing using standard microscope control hardware—such as a desktop computer—with sufficient performance. HPC workflows do not only come with large costs and limited accessibility, but they potentially also have higher latency than local computing infrastructure, stemming from e.g. job scheduling and data transfer to and from shared storage. Such high latency is an important limitation for HPC when aiming for truly live processing, for example in using reconstructions from data streams in a search mode when finding a sample or optimizing the optics of the microscope. This is an important consideration as, for example, the benefits provided by dose-efficient 4D-STEM methods are strongly mitigated when the user is still limited to less appropriate methods like ADF imaging when setting up for acquisition, and is forced to choose between either damaging the sample anyway or working blindly. These considerations regarding the implications of HPC for 4D-STEM data processing, like cost, barrier to entry, accessibility and latency, are important to consider now, as they determine the future of these 4D-STEM methods and their integration in the wider field. Aiming for accessible, fast and efficient workflows increases the chances of 4D-STEM methods to establish themselves as routine and widespread tools for the investigation of e.g. life science specimens or other beam-sensitive materials.
@@
In the discussion above, the distinction between frame-based and event-driven detectors is made on the level of the chip application-specific integrated circuit (ASIC), as in the case of the Medipix3 and Timepix3 chips. However, the read-out mechanism of a DED does not necessarily have to be considered fixed to that of the ASIC, as it is possible to transform the initial data stream between different representations [53, 54, 83, 12]. This transformation could happen on a low hardware level, before the data stream reaches the user’s processing computer, for example by performing run-length encoding using a field-programmable gate array (FPGA). Many of the properties of the DED will remain determined by the configuration of the ASIC, like the maximum beam current, the maximum speed, and the information content, e.g. the presence of TOA information. Consequently, this will determine the range of applications allowed by the DED, such as e.g. time-resolved measurements [46, 47, 43, 45, 44, 48].The arguments presented above remain valid irrespective of the initial configuration of the DED, as these simply consider sparse representations of electron data rather than the more classical dense representation.
@@
Currently, the event-driven DED for which acquisition control is already implemented include the Amsterdam Scientific Instruments CheeTah T3 and a custom prototype built from a commercial Advacam AdvaPIX T3. The frame-based Quantum Detectors MerlinEM [91, 92] is also supported, which offers acquisition in a 1 bit mode, having interests for straightforward sparsification [40]. Regarding scan control, the framework currently supports the use of the Quantum Detectors scan engine, which provides possibilities for various triggering and synchronization functionalities including e.g. line triggering, pixel triggering, master clock generation and pattern scanning. These functionalities are aligned with the processing pipeline to enable seamless synchronization in many configurations. For the integration of the microscope column’s hardware and functionalities, there currently is support for Thermo Fisher Scientific instruments employing the temscript Python wrapper [93]. Integration of JEOL systems is planned for future work. A graphical user interface (GUI) is available providing an interface to the centralized and synchronized control over all hardware components and live feedback of experimental output.
@@
Live processing of the raw data streams provided by the detector, encompassing e.g. the decoding of raw events to usable electron coordinates, requires processing rates of the order of GB/s. These performance requirements warrant the need for an efficient implementation in a high-performance language, which is in our case the C++ language. Another framework for the processing of event-driven STEM experiments, dedicated towards EELS, was developed by Auad et al. [50] and is based on the Rust programming language.
@@
In the case of the event-driven kernel-additive approach, the computational complexity of the reconstruction follows $\text{O}(N\cdot R^{2})$, with $N$ the number of events and $R^{2}$ the total amount of pixels in the kernel. As such, the number of floating point operations per second (FLOPS) that has to be made to live process a complete event-driven data stream, in the GPRI framework, can be directly estimated from the beam current and kernel size. A single-chip Timepix3 detector may provide a maximum of about 80 million events per second, which corresponds to a beam current of a few pA depending on the cluster size. A realistic size for a GPRI kernel is 64 $\times$ 64 pixels in the reconstruction window, and a single-precision floating-point representation is sufficient. In this scenario of a single precision 64 $\times$ 64 kernel, and assuming a beam current of 5 pA with an average cluster size of 3 pixels, the number of additions to be performed per second is just under 400 GFLOP. Going further, a more detailed evaluation of the required FLOPS, as a function of beam current or frame rate and kernel size, is presented in figure 5. Standard graphical processing units (GPU) available in desktop computers, like the Nvidia RTX series, have a theoretical peak performance in the tens to hundreds of TFLOPS regime. While the FLOPS performance achieved in practice is quite significantly less than this theoretical value, it then remains clear that the computational requirements are within an achievable order of magnitude. More generally, this means that, in the GPRI framework, there is no requirement for advanced hardware, that would go beyond standard consumer hardware, to meet the computational requirements of live analytical ptychography reconstructions from DED data streams. This nevertheless still requires a performant implementation of the processing pipeline, that can sufficiently leverage the parallel nature of a GPU.
Figure 5: Assessment of the computational requirements of a kernel-additive method like GPRI as a function of kernel size and data rate. In a, the event hit rate and associated beam current (with a cluster size 3 pixels) is used, showing the maximum hit rate of current generation event-driven DED. In b, the frame rate and associated frame time is used, situating current generation DED [9, 91, 11]. To enable the comparison in b, a detector of 128 $\times$ 128 pixels is assumed, which is not the actual total number of pixels for the detectors marked with a star. The displayed FLOPS range is fixed to facilitate the comparison between a and b, with markings indicating the minimum and maximum value of each. The required FLOPS range from the order of GFLOPS to 100 TFLOPS, which is within the theoretical peak performance of a single current-generation consumer GPU.
@@
In this subsection, a demonstration is made of the live processing performance using a consumer desktop computer consisting of a 16 core AMD Ryzen threadripper pro 5995wx CPU and an Nvidia RTX4090 GPU. Although this is a high-end desktop configuration, it remains extremely modest in comparison to typical HPC infrastructure. For repeatability, the evaluations are based on raw recorded datasets stored on an NVMe solid-state disk. However, the results are also representative for raw data streams coming from a socket connected to the detector. For GPRI, GPU acceleration is achieved using a combination of Pytorch’s LibTorch API [98] and custom CUDA kernels [99]. Before making an assessment of the processing speeds of the ptychographic reconstructions using the GPRI framework, an evaluation is made of the performance of the underlying evenTem pipeline for parsing the raw packets to $\left(r_{x},r_{y},k_{x},k_{y}\right)$ coordinates. To this end, a virtual ADF reconstruction is made using a single CPU thread for reading data and a single CPU thread for parsing the events. The riCoM method is also used for comparison, as it has already been shown to provide live dose-efficient reconstructions, based directly on the event stream.
@@ Discussion
  • When the time resolution of the detector is in the nanosecond or better regime, as in the case of Timepix3 and Timepix4, the detector does not impose the practical limit on the scan speed. Instead, this limit is imposed by response time of the scan system [49, 100].
@@
Overall, this means that Timepix4-based detectors hold promises to become an ideal all-around solution for STEM, thus reinforcing the need to drive the integration of event-driven operation into electron microscopy workflows. The first experimental setups integrating the Timepix4 chip have already been realized [59], although the maximum data rate performance has not yet been reported. From the assessments presented in this work, it is clear that the event format still provides numerous benefits over a dense format in many of the scenarios enabled by Timepix4. When utilizing its full potential bandwidth, Timepix4 would allow event rates far exceeding those generated by e.g. a typical 50 pA electron beam. To keep up with such an increased data rate, the presented processing pipeline will have to allow further parallelization. The current workflow uses a single CPU thread for data reading and a single CPU thread for event parsing. If events can be parsed independently of each other, an expansion to multiple CPU threads or a GPU should be straightforward. As discussed above, when synchronization is ensured by the use of a master clock, events can indeed be treated independently. A tenfold increase of the current processing rate, e.g. achieved through parallelization, would already allow live processing of the data stream resulting from a typical 50 pA electron beam.